OpenStudy (anonymous):

The distance between the center of symmetry of a parallelogram and its longer side is equal to 12 cm. The area of the parallelogram is 720 cm^2 and its perimeter is 140 cm. Determine the length of the longer diagonal of the parallelogram. please help :(

hartnn (hartnn):

Let the sides be 'x' and 'y' Perimeter = 2 (x+y) =14 0 so, x+y =70 Area = xy = 720 can you find the sides from here first ? x, y ?

OpenStudy (anonymous):

i am not really sure how, sorry :(

hartnn (hartnn):

the centre of symmetry is the intersection of diagonals, right ?? ok, can you verify the question once please ? because i am getting irrational values of 'x' and 'y' .....they may be correct, but still...

OpenStudy (anonymous):

 if my drawing is right, then it means that x = 2(12) right? then 140= 2(24)+2(y)... y = 46 and x = 24 am i right?

hartnn (hartnn):

since its the distance given, i assume it will be  but if its what you have done, then it simplify things...